OpenStudy (anonymous):

okay, so i have a problem which i have solved to a point, but i was wondering if someone could help clarify the steps for me? it is : cosx+1=1, and i changed it so it became cosx=0, but i'm not sure if i'm doing it correctly or if my steps are correct.Thanks!

OpenStudy (anonymous):

cosx + 1= 1 - 1 -1 ........ Subtract 1 from both sides cox x = 0

OpenStudy (whpalmer4):

Yes, that's a reasonable step, though we can't guarantee that what came before is correct...

OpenStudy (anonymous):

all that was given was cosx+1=0, with the interval [0,2pi), and to find the solutions. I reached this point, but I'm not sure where to go from there. :) Thanks for helping.

OpenStudy (whpalmer4):

Wait, \[\cos x + 1 = 0\] or \[\cos x + 1 = 1\]?

OpenStudy (anonymous):

whoops, my bad. it is supposed to be cosx+1=1, i should have been more careful,

OpenStudy (anonymous):

but the rest is acccurate

OpenStudy (phi):

do you know what the plot of cos(x) looks like between 0 and 2 pi ?

OpenStudy (anonymous):

I know that it makes a complete round of the unit circle, right?

OpenStudy (phi):

see http://www.mathsisfun.com/algebra/trig-sin-cos-tan-graphs.html scroll down to cos graph

OpenStudy (anonymous):

I see... so it is this one right http://www.mathsisfun.com/algebra/images/cosine-graph.gif

OpenStudy (phi):

yes. now list the x's where cos(x) is 0

OpenStudy (anonymous):

I'm not sure, is it 1 and -1?

OpenStudy (phi):

put your finger on the "blue line" (the cosine curve) and trace it until you reach the x-axis. then look down to see what x value that is (between 0 and 2pi)

OpenStudy (anonymous):

oh, i see, it's -3pi/2,-pi/2, pi/2,3pi/2

OpenStudy (anonymous):

thank you

OpenStudy (phi):

yes, but they only want the values in the interval [0,2pi) the bracket [0 means from 0 the "close parens" 2 pi ) means up to but not including 2 pi in this case, there are only the 2 values pi/2 and 3pi/2

OpenStudy (anonymous):

These values are familiar, so do they both count as a solution?

OpenStudy (phi):

Familiar or not, they are the two places in the interval [0, 2pi) where cos(x) is 0

OpenStudy (anonymous):

okay, once the problem is solved, by using the cos graph to see where the values match up, thus giving the correct solution

OpenStudy (anonymous):

Thanks everyone! This really helped! Have a good day!

OpenStudy (phi):

yes. But most people memorize the sin, cos and tan of a few special angles.

OpenStudy (anonymous):

That makes sense. If you don't mind my asking, what angles are these exactly, just so i know what to keep track of? If not, that's okay, you've helped me a lot already. :)

OpenStudy (phi):

The angles (in degrees) are 0, 30, 45, 60, and 90 (in radians 0, pi/6, pi/3, pi/4, 2pi/3 and pi/2 ) angle sin cos tan 0 0 1 0 30 ½ sqr(3)/2 1/sqr(3) 45 sqr(2)/2 sqr(2)/2 1 60 sqr(3)/2 ½ sqr(3) 90 1 0 not defined

OpenStudy (anonymous):

This is so helpful! Thank you for your time and effort!